# Gaussian Processes for Time Series with Lead-Lag Effects with application to biology data

### v1

@@ 5 Real-world application: Dynamic chromatin interactions
Secondly, we analyzed the enrichment of looped enhancer-gene pairs between the two groups. A looped enhancer-gene pair is one where the enhancer element is physically brought into close proximity with the regulated gene, enabling the enhancer’s activity to influence the expression of the associated gene. Looping in this dataset was determined using Hi-C counts, but the determination of loops uses spatial pattern detection and statistical models to determine if the signal rises to the level of a “loop”. We observed that candidate enhancer-gene pairs identified by GPlag had a significantly higher likelihood of being ”looped” enhancer-gene pairs compared to the matched pairs from the other group, selected based on genomic distance (Davis et al., 2022) ($\chi^{2}$ test, $p=0.05$). However, no significant difference in enrichment was found between the two groups of pairs identified using baseline methods.
@@
[…]
Figure 5: Visualization of GPlag modelling output on selected pairs. The title for each subfigure represents the enhancer’s genomic location and the gene name. The grey line represents gene expression; the blue line represents enhancer activity; the red line represents the enhancer shifted by GPlag; and the brown line represents the enhancers shifted by TLCC.
@@ 6 Discussion
The work here suggests some exciting future directions. First, when the sample size $n$ is small, the choice of the kernel is crucial in GP models. The extension to a non-stationary kernel allows for more flexibility when modelling non-stationary time series with more heterogeneity. The two main challenges are to find a valid non-stationary kernel with interpretable parameters and perform efficient and effective statistical inference. On the other hand, for time series data with large $n$, i.e., dense time points, the vanilla GP is computationally expensive with $\mathrm{O}(n^{3})$ complexity (Rasmussen, 2004). In this situation, we can apply scalable GP methods such as nearest neighbor Gaussian processes (NNGP, Datta et al. (2016)) or variational inference GP (Tran et al., 2015) to reduce the complexity to $\mathrm{O}(n\log n)$ for large datasets.
Appendix A Additional covariance functions
@@
$K((t,l),(t^{\prime},l^{\prime}))$ $=\frac{2\sigma^{2}c^{1/2}\left\{\frac{b}{2}\left(\frac{a_{ll^{\prime}}^{2}+1}{a_{ll^{\prime}}^{2}+c}\right)^{1/2}|t-t^{\prime}+s_{l}-s_{l^{\prime}}|\right\}^{\nu}}{(a_{ll^{\prime}}^{2}+1)^{\nu}(a_{ll^{\prime}}^{2}+c)^{1/2}\Gamma(\nu)}K_{\nu}\left(b\left(\frac{a_{ll^{\prime}}^{2}+1}{a_{ll^{\prime}}^{2}+c}\right)^{1/2}|t-t^{\prime}+s_{l}-s_{l^{\prime}}|\right),$ (6)
$K((t,l),(t^{\prime},l^{\prime}))$ $=\frac{\sigma^{2}c^{1/2}}{(a_{ll^{\prime}}^{2}+1)^{1/2}(a_{ll^{\prime}}^{2}+c)^{1/2}}\exp\left\{-b\left(\frac{a_{ll^{\prime}}^{2}+1}{a_{ll^{\prime}}^{2}+c}\right)^{1/2}|t-t^{\prime}+s_{l}-s_{l^{\prime}}|\right\},$ (7)
$K((t,l),(t^{\prime},l^{\prime}))$ $=\frac{\sigma^{2}}{(a_{ll^{\prime}}^{2}+1)^{1/2}}\exp\left\{-\frac{b|t-t^{\prime}+s_{l}-s_{l^{\prime}}|}{(a_{ll^{\prime}}^{2}+1)^{1/2}}\right\},$ (8)
$K((t,l),(t^{\prime},l^{\prime}))$ $=\frac{\sigma^{2}(a_{ll^{\prime}}^{2}+1)^{1/2}}{(a_{ll^{\prime}}^{2}+1)+b(t-t^{\prime}+s_{l}-s_{l^{\prime}})^{2}},$ (9)
$K((t,l),(t^{\prime},l^{\prime}))$ $=\sigma^{2}\exp\left\{-a_{ll^{\prime}}^{2}-b(t-t^{\prime}+s_{l}-s_{l^{\prime}})^{2}-a_{ll^{\prime}}^{2}(t-t^{\prime}+s_{l}-s_{l^{\prime}})^{2}\right\},$ (10)
As an illustration, Equation 6 is the product of a kernel in Gneiting family with $\phi(t)=\frac{(\tilde{b}t^{1/2})^{\nu}}{2^{\nu-1}\Gamma(\nu)}K_{\nu}(\tilde{b}t^{1/2})$, $\psi(t)=(a^{2}t+c)/c(a^{2}t+1)$ in Corollary 1 with $\tilde{b}=b/\sqrt{c}$, and a simple kernel $\frac{1}{(a_{ll^{\prime}}^{2}+1)^{\nu+1/2}}$; Equation 7 is a special case of Equation 6 when $\nu=1/2$. Moreover, if we set $c=1$, it becomes LExp; Equation 8 can be derived by setting $\phi(t)=e^{-bt^{1/2}}$, $\psi(t)=a^{2}t+1$. Equation 9 can be derived by setting $\phi(t)=(1+bt)^{-1}$, $\psi=a^{2}t+1$. Equation 10 is not in Gneiting family, but from Cressie and Huang (1999), who proposed another spectral density based construction for nonseparable spatiotemporal kernels.
@@
Appendix B Proof of Theorem 1 and Theorem 2
@@
$K((t,l),(t^{\prime},l^{\prime}))\coloneqq K^{\prime}(t-t^{\prime}+s_{l}-s_{l^{\prime}},l,l^{\prime})$ (11)
@@
where $\Sigma^{\prime}$ is the covariance matrix derived from $K^{\prime}$, a positive definite kernel, evaluated on the “fake” set of observations, hence is positive definite. As a result, $K$ is positive definite for any time lags $\{s_{i}\}_{l=1}^{L}$.
@@ Appendix C Proof of Theorem 3
The spectral densities of LRBF, LMat and LExp are given by
@@
Now we can prove Theorem 3. Observe that
@@
$e^{i\omega(s-\widetilde{s})}\frac{\alpha b^{2\nu}}{\widetilde{\alpha}\widetilde{b}^{2\nu}}=1.$ (12)
First we let $l=l^{\prime}$ so that $s=s_{l}-s_{l^{\prime}}=0$ and $a_{ll^{\prime}}=0$. As a result, Equation (12) becomes
$\sigma^{2}b^{2\nu}=\widetilde{\sigma}^{2}\widetilde{b}^{2\nu}.$ (13)
Then let $\omega=0$, and plug Equation (13) into Equation (12), we have
$\frac{1}{a_{ll^{\prime}}^{2}+1}=\frac{1}{\widetilde{a}_{ll^{\prime}}^{2}+1},\leavevmode\nobreak\ \text{i.e.}\leavevmode\nobreak\ a_{ll^{\prime}}=\widetilde{a}_{ll^{\prime}}.$ (14)
Finally, plug Equation (13) and (14) into Equation (12), we have $e^{i\omega(s-\widetilde{s})}=1$, that is, $s_{l}-s_{l^{\prime}}=\widetilde{s}_{l}-\widetilde{s}_{l^{\prime}}$. Recall that $s_{1}=\widetilde{s}_{1}=0$ by the definition, so we let $l^{\prime}=1$ to get
$s_{l}=\widetilde{s}_{l},\leavevmode\nobreak\ l=1,\cdots,L.$ (15)
Combination Equations (13), (14), and (15), we conclude that
@@
To show finite nature of the integral necessitating the condition of (12), we assume $L=2$ and $\nu=1/2$, i.e., there are only two time series with LExp kernel. In this situation, it suffices to show that if $(\sigma^{2}b,a,s)\neq(\tilde{\sigma}^{2}\tilde{b},\tilde{a},\tilde{s})$, then $K(\cdot;\sigma^{2},b,a,s)\not\equiv K(\cdot;\tilde{\sigma}^{2},\tilde{b},\tilde{a},\tilde{s})$.
@@
Similarly, when $a\neq\tilde{a}$, we can prove the same result with $\eta=(\tilde{a}^{2}+1)/(a^{2}+1)$; while when $s\neq\tilde{s}$, the corresponding $\eta=e^{-b\tilde{s}}/e^{-bs}$.
Appendix D Supplementary Algorithm
@@
  Step 3: Maximize the above log-likelihood by user selected optimizer such as Adam or L-BFGS-B.
@@
  Step 3: Sampling: Run No-U-Turn Sampler(NUTS) algorithm to generate samples from the above posterior distribution to achieve estimator uncertainty.
@@
Appendix E Additional experimental details
E.1 Simulation
Parameter Estimation and Inference The variable $T$ with various sample size was generated within the range (-50, 50), and a random variation was added by sampling from a uniform distribution with a range of $\mathrm{unif}(-\frac{1}{4},\frac{1}{4})$. Three kernels were all served as a covariance function to generate pairwise time series from Gaussian process(GP) with $b$, $a$, $s$ set to $0.3,1,2$. When fitting GPlag, the parameters $b$, $a$, $s$ were initialized with 1, 1, 1, respectively. For both GPlag and Lead-Lag methods, the upper bound for the time lag was set to 4. The maximum likelihood estimation (MLE) was obtained using the L-BFGS-B algorithm.
@@ E.1 Simulation
Figure 6: Prediction accuracy comparing with baselines for pair time series generated from a linear process $y=2t+3+5\epsilon_{1}$. $\epsilon_{1}$ was sampled from t-distributed noise. Each time series has $n=100$ observations, whereas the time shift =20 between time series. Replicates 100 times.
Clustering or Ranking Time Series based on dissimilarity to target time series The variable $T$ with $n_{l}=50$ for 10 time series (target time series + other 9 time series) was generated regularly from (-2, 2). When fitting GPlag, the parameters $b$, $a$, $s$ were initialized with 1, 1, 0, respectively. The lower and upper bound for the time lag was set to -1 and 4, respectively. The maximum likelihood estimation (MLE) was obtained using the L-BFGS-B algorithm.
E.2 Real-world applications
[…]

### v2

@@ 5 Genomics application: Dynamic chromatin interactions
Secondly, we analyzed the enrichment of looped enhancer-gene pairs between the two groups. A looped enhancer-gene pair is one where the enhancer element is physically brought into close proximity with the regulated gene, enabling the enhancer’s activity to influence the expression of the associated gene. Looping in this dataset was determined using Hi-C counts, but the determination of loops uses spatial pattern detection and statistical models to determine if the signal rises to the level of a “loop”. We observed that candidate enhancer-gene pairs identified by GPlag had a significantly higher likelihood of being “looped” enhancer-gene pairs compared to the matched pairs from the other group, selected based on genomic distance (Davis et al., 2022) ($\chi^{2}$ test, $p=0.05$). However, no significant difference in enrichment was found between the two groups of pairs identified using baseline methods.
@@
Thirdly, to evaluate if our candidate enhancer-promoter pairs were detected as having a functional relationship in an orthogonal experimental data type, we downloaded the genomic locations of response expression quantitative trait loci (response eQTL) found in human macrophages exposed to IFNγ (Alasoo et al., 2018) (that is, using one of the two stimuli used in the Reed et al. (2022) experiment, and in the same cell type). Because the number of eQTLs was small compared to the number of enhancers, we used a relatively stringent threshold to label a putative functional enhancer-promoter pair: we split the candidate enhancer-promoter pairs into groups according to posterior mean correlation ($>0.98$). There were 74 H3K27ac peaks overlapping with macrophage response eQTL, among which 12 pairs were in our defined candidate enhancer-promoter pair sets. Therefore, there was a significant enrichment of our candidate enhancer-promoter pairs to eQTLs ($\chi^{2}$ test $p=0.18\times 10^{-3}$) futher suggesting the relevance of the enhancer-promoter pairs highly ranked by GPlag.
Figure 5: Visualization of GPlag modelling output on selected pairs. The title for each subfigure represents the enhancer’s genomic location and the gene name. The solid circle line represents gene expression; the solid triangle line represents enhancer activity; the dashed circle line represents the enhancer shifted by GPlag; and the dashed triangle line represents the enhancers shifted by TLCC.
@@ 6 Discussion
The work here suggests some exciting future directions. First, when the sample size $n$ is small, the choice of the kernel is crucial in GP models. The extension to a non-stationary kernel allows for more flexibility when modelling non-stationary time series with more heterogeneity. The two main challenges are to find a valid non-stationary kernel with interpretable parameters and perform efficient and effective statistical inference. On the other hand, for time series data with large $n$, i.e., dense time points, the vanilla GP is computationally expensive with $\mathrm{O}(n^{3})$ complexity (Rasmussen, 2004). In this situation, we can apply scalable GP methods such as nearest neighbor Gaussian processes (NNGP, Datta et al. (2016)) or variational inference GP (Tran et al., 2015) to reduce the complexity to $\mathrm{O}(n\log n)$ for large datasets. In addition, incorporating additional parameters, such as varying the variance $\sigma^{2}$ and lengthscale $b$ across different time series, could potentially enhance model performance. Finally, the question of whether the parameters in Theorem 3 can be consistently estimated remains open. Even for simpler kernels such as the Matérn, consistent estimators for identifiable parameters (e.g., $\sigma^{2}b^{2\nu}$ and $\nu$) were only recently developed (Loh et al., 2021; Loh and Sun, 2023). For more complex kernels, like the ones proposed in this paper, constructing consistent estimators is an open and challenging problem.
Appendix A Additional covariance functions for GPlag
@@
$K((t,l),(t^{\prime},l^{\prime}))$ $=\frac{2\sigma^{2}c^{1/2}\left\{\frac{b}{2}\left(\frac{a_{ll^{\prime}}^{2}+1}{a_{ll^{\prime}}^{2}+c}\right)^{1/2}|t-t^{\prime}+s_{l}-s_{l^{\prime}}|\right\}^{\nu}}{(a_{ll^{\prime}}^{2}+1)^{\nu}(a_{ll^{\prime}}^{2}+c)^{1/2}\Gamma(\nu)}K_{\nu}\left(b\left(\frac{a_{ll^{\prime}}^{2}+1}{a_{ll^{\prime}}^{2}+c}\right)^{1/2}|t-t^{\prime}+s_{l}-s_{l^{\prime}}|\right),$ (S1)
$K((t,l),(t^{\prime},l^{\prime}))$ $=\frac{\sigma^{2}c^{1/2}}{(a_{ll^{\prime}}^{2}+1)^{1/2}(a_{ll^{\prime}}^{2}+c)^{1/2}}\exp\left\{-b\left(\frac{a_{ll^{\prime}}^{2}+1}{a_{ll^{\prime}}^{2}+c}\right)^{1/2}|t-t^{\prime}+s_{l}-s_{l^{\prime}}|\right\},$ (S2)
$K((t,l),(t^{\prime},l^{\prime}))$ $=\frac{\sigma^{2}}{(a_{ll^{\prime}}^{2}+1)^{1/2}}\exp\left\{-\frac{b|t-t^{\prime}+s_{l}-s_{l^{\prime}}|}{(a_{ll^{\prime}}^{2}+1)^{1/2}}\right\},$ (S3)
$K((t,l),(t^{\prime},l^{\prime}))$ $=\frac{\sigma^{2}(a_{ll^{\prime}}^{2}+1)^{1/2}}{(a_{ll^{\prime}}^{2}+1)+b(t-t^{\prime}+s_{l}-s_{l^{\prime}})^{2}},$ (S4)
$K((t,l),(t^{\prime},l^{\prime}))$ $=\sigma^{2}\exp\left\{-a_{ll^{\prime}}^{2}-b(t-t^{\prime}+s_{l}-s_{l^{\prime}})^{2}-a_{ll^{\prime}}^{2}(t-t^{\prime}+s_{l}-s_{l^{\prime}})^{2}\right\},$ (S5)
As an illustration, Equation S1 is the product of a kernel in Gneiting family with $\phi(t)=\frac{(\tilde{b}t^{1/2})^{\nu}}{2^{\nu-1}\Gamma(\nu)}K_{\nu}(\tilde{b}t^{1/2})$, $\psi(t)=(a^{2}t+c)/c(a^{2}t+1)$ in Corollary 1 with $\tilde{b}=b/\sqrt{c}$, and a simple kernel $\frac{1}{(a_{ll^{\prime}}^{2}+1)^{\nu+1/2}}$; Equation S2 is a special case of Equation S1 when $\nu=1/2$. Moreover, if we set $c=1$, it becomes LExp; Equation S3 can be derived by setting $\phi(t)=e^{-bt^{1/2}}$, $\psi(t)=a^{2}t+1$. Equation S4 can be derived by setting $\phi(t)=(1+bt)^{-1}$, $\psi=a^{2}t+1$. Equation S5 is not in Gneiting family, but from Cressie and Huang (1999), who proposed another spectral density based construction for nonseparable spatiotemporal kernels.
@@
A.1 Proof of Theorem 1 and 2
@@
$K((t,l),(t^{\prime},l^{\prime}))\coloneqq K^{\prime}(t-t^{\prime}+s_{l}-s_{l^{\prime}},l,l^{\prime})$ (S6)
@@
where $\Sigma^{\prime}$ is the covariance matrix derived from $K^{\prime}$, a positive definite kernel, evaluated on the “fake” set of observations, hence is positive definite. As a result, $K$ is positive definite for any time lags $\{s_{i}\}_{l=1}^{L}$. ∎
@@ A.2 Proof of Theorem 3
The spectral densities of LRBF, LExp and LMat and are given by
@@
Now we can prove Theorem 3.
Proof of Theorem 3.
Observe that
@@
$e^{i\omega(s-\widetilde{s})}\frac{\alpha b^{2\nu}}{\widetilde{\alpha}\widetilde{b}^{2\nu}}=1.$ (S7)
Since Equation (S7) holds for any $l,l^{\prime},\omega$, we set certain parameters to special values to isolate conditions on certain parameters. First we let $l=l^{\prime}$ so that $s=s_{l}-s_{l^{\prime}}=0$ and $a_{ll^{\prime}}=0$. As a result, Equation (S7) becomes
$\sigma^{2}b^{2\nu}=\widetilde{\sigma}^{2}\widetilde{b}^{2\nu}.$ (S8)
Then let $\omega=0$, and plug Equation (S8) into Equation (S7), we have
$\frac{1}{a_{ll^{\prime}}^{2}+1}=\frac{1}{\widetilde{a}_{ll^{\prime}}^{2}+1},\leavevmode\nobreak\ \text{i.e.}\leavevmode\nobreak\ a_{ll^{\prime}}=\widetilde{a}_{ll^{\prime}}.$ (S9)
Finally, plug Equation (S8) and (S9) into Equation (S7), we have $e^{i\omega(s-\widetilde{s})}=1$, that is, $s_{l}-s_{l^{\prime}}=\widetilde{s}_{l}-\widetilde{s}_{l^{\prime}}$. Recall that $s_{1}=\widetilde{s}_{1}=0$ by the definition, so we let $l^{\prime}=1$ to get
$s_{l}=\widetilde{s}_{l},\leavevmode\nobreak\ l=1,\cdots,L.$ (S10)
Combination Equations (S8), (S9), and (S10), we conclude that
@@
To show finite nature of the integral necessitating the condition of (S7), we assume $L=2$ and $\nu=1/2$, i.e., there are only two time series with LExp kernel. In this situation, it suffices to show that if $(\sigma^{2}b,a,s)\neq(\tilde{\sigma}^{2}\tilde{b},\tilde{a},\tilde{s})$, then $K(\cdot;\sigma^{2},b,a,s)\not\equiv K(\cdot;\tilde{\sigma}^{2},\tilde{b},\tilde{a},\tilde{s})$.
@@
Similarly, when $a\neq\tilde{a}$, we can prove the same result with $\eta=(\tilde{a}^{2}+1)/(a^{2}+1)$; while when $s\neq\tilde{s}$, the corresponding $\eta=e^{-b\tilde{s}}/e^{-bs}$. ∎
A.3 Detailed algorithm
@@
1  Step 3: Maximize the above log-likelihood by user selected optimizer such as Adam or L-BFGS-B.
@@
1  Step 3: Sampling: Run No-U-Turn Sampler(NUTS) algorithm to generate samples from the above posterior distribution to achieve estimator uncertainty.
@@
Appendix B Additional experimental details
B.1 Simulation
Parameter Estimation and Inference The variable $T$ with various sample size was generated within the range (-50, 50), and a random variation was added by sampling from a uniform distribution with a range of $\mathrm{unif}(-\frac{1}{4},\frac{1}{4})$. Three kernels were all served as a covariance function to generate pairwise time series from Gaussian process(GP) with $b$, $a$, $s$ set to $0.3,1,2$. When fitting GPlag, the parameters $b$, $a$, $s$ were initialized with 1, 1, 1, respectively. For both GPlag and Lead-Lag methods, the upper bound for the time lag was set to 4. The maximum likelihood estimation (MLE) was obtained using the L-BFGS-B algorithm. We plot the MLE of $b$ and $\sigma^{2}$, as well as $\sigma^{2}b$ (Figure S1). The value of $\sigma^{2}b$ closely matches its true value, aligning with the theoretical predictions discussed in Theorem Theorem 3. Additionally, as $n$ increases, the variance decreases, demonstrating the consistency of the MLE.
Furthermore, to evaluate the method’s performance across different lengthscales $b$, and to examine their influence on the estimates of $a$ and $s$, we simulate $b=0.3,5,10$ for $n=50,100,200,300$ using the LExp kernel. The results, summarized in the Figure S2, indicate that the value of $b$ does influence model optimization. Nonetheless, MLE of both $a$ and $a$ converge to their true values as the sample size increases.
Figure S1: MLEs of GPlag parameters $b$ and $\sigma^{2}$ and $\sigma^{2}b$ in the LExp kernel with different n. The true parameter values were represented by the dashed red horizontal lines. All boxplots were from 100 replicates
Figure S2: MLE of $a$ and $s$ in GPlag with data generated from LExp kernel with different lengthscales $b$ and sample size. The true parameter values were represented by the dashed red horizontal lines. All boxplots were from $20$ replicates
@@ B.1 Simulation
Figure S3: Prediction accuracy comparing with baselines for pair time series generated from a linear process $y=2t+3+5\epsilon_{1}$. $\epsilon_{1}$ was sampled from t-distributed noise. Each time series has $n=100$ observations, whereas the time shift =20 between time series. Replicates 100 times.
Clustering or Ranking Time Series based on dissimilarity to target time series The variable $T$ with $n_{l}=50$ for 10 time series (target time series + other 9 time series) was generated regularly from (-2, 2). When fitting GPlag, the parameters $b$, $a$, $s$ were initialized with 1, 1, 0, respectively. The lower and upper bound for the time lag was set to -1 and 4, respectively. The maximum likelihood estimation (MLE) was obtained using the L-BFGS-B algorithm. We also replicated this experiment using LRBF and LMat kernels. For both kernels, the ARI and NMI scores were 1 (same as LExp kernel), and the inferred values of $a$ were again clearly separated into three distinct clusters Figure S4.
Figure S4: Inferred $a$ with LRBF and LMat kernels in the clustering experiment.
Visual example of kernels
To better help the reader understand and choose the appropriate kernel, here we simulate data from LExp, LMat, and LRBF with various parameter. The simulated sample paths are shown in the following figures. The figures clearly demonstrate that a smaller $b$ indicates higher similarity between nearby points, while a smaller $a$ reflects greater similarity between the two time series.
Figure S5: Simulated sample path for LExp kernel.
Figure S6: Simulated sample path for LRBF kernel.
Figure S7: Simulated sample path for LMat ($\nu=\frac{3}{2}$) kernel.
B.2 Real-world applications
In order to understand how enhancers regulate gene transcription, Reed et al. (2022) evaluated human macrophage activation after stimulating macrophage with LPS and interferon-gamma (IFNγ). The data were collected at seven irregularly-spaced time points (0, 0.5, 1, 1.5, 2, 4, 6 h). At each time point, the 3D chromatin structure was profiled using in situ Hi-C, putative enhancer activity using chromatin immunoprecipitation sequencing (ChIP-seq) targeting histone 3 lysine 27 acetylation (H3K27ac), chromatin accessibility using assay for transposase-accessible chromatin using sequencing (ATAC-seq), and gene expression using RNA sequencing (RNA-seq). Hi-C scale was indicated in KR-normalized counts, H3K27ac and RNA-seq using variance stabilizing transformation (VST) from the DESeq2 package (Love et al., 2014).